OpenStudy (anonymous):

deterime whether the sequnce is geometric 9,17,25,33... 4,12,36,108... -320,-80,-20,-5... -75,-60,-45,-30... 64,-32,16,-8... 26.5,21,15.5,10...

OpenStudy (anonymous):

You take the 2nd term divided by the 1st term, the 3rd term divided by the 2nd term, the 4th term divided by the 3rd term. I did #1 for you. Is 17/9 = 25/17 = 33/25 if they are all the same number then it IS geometric, if they are NOT ALL the same them it IS NOT geometric.

OpenStudy (anonymous):

thanks i fell asleep during this lesson lol